# Mixture-of-Skills: Learning to Optimize Data Usage for Fine-Tuning Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated their extraordinary capabilities and are expected to proficiently master a diverse range of skills (Ouyang et al., 2022; Sanh et al., 2022; OpenAI, 2023; Anil et al., 2023b; Touvron et al., 2023a, b; Anil et al., 2023a; Mesnard et al., 2024), such as writing, reasoning, chatting, coding, and more, through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) on an extensive collection of datasets from various sources (Bai et al., 2022; Longpre et al., 2023; Ding et al., 2023). Each dataset contributes unique elements to the model’s skill set, but this diversity also brings challenges.
@@ 2 Preliminaries
Wang et al. (2020a) propose a general framework that automatically re-weighs training instances to enhance model performance, utilizing a validation set $D_{\textrm{dev}}$. This framework consists of two components: the model $\boldsymbol{\theta}$ and the scorer network $\boldsymbol{\psi}$. The scorer network $\boldsymbol{\psi}$ is designed to calculate a sampling probability for each training instance, which reflects its impact on validation performance. Training instances that have a greater similarity with $D_{\textrm{dev}}$ are allocated a higher probability, thus increasing their likelihood of being selected for model $\boldsymbol{\theta}$ updates. MoS is inspired by DDS but has key differences. Firstly, MoS focuses on rebalancing datasets, unlike DDS, which reweighs training instances. Secondly, MoS does not require prior knowledge of downstream applications, whereas DDS relies on validation set feedback, risking overfitting to that specific validation set. Thirdly, DDS uses the same architecture for the scorer network $\boldsymbol{\psi}$ and model $\boldsymbol{\theta}$, limiting its scalability, while MoS uses a simple MLP model as its scorer network.
@@ 3.2.3 Learning Trajectory
4 Experiments
4.1 Experimental Setup
Datasets
In this work, we collect supervised fine-tuning (SFT) datasets for four distinct skills: Mathematics (Yue et al., 2023), Medicine (Zhang et al., 2023a), General (ShareGPT), [^1] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered and NLP (Sanh et al., 2022). Due to the constraint on compute, we sample 10% of examples from each dataset. More details can be found in Appendix B.
@@ 4.1 Experimental Setup
We apply MoS to three diverse model backbones, including Qwen1.5-0.5B (Bai et al., 2023), Gemma-2B (Mesnard et al., 2024), and Llama-3-8B (AI@Meta, 2024). Optimization details are in Appendix C.
@@
We compare MoS with three heuristic baselines based on Equation 3: proportional sampling (Prop., $\tau=1$), temperature sampling (Temp., $\tau=10$), and uniform sampling (Uni., $\tau=\infty$). We do not use the maximum cap and other instance selection methods as our baselines because they fail to fully utilize all available data.
@@
In this work, we evaluate our models on two widely-used benchmarks that are highly correlated with human judgments: MMLU (Hendrycks et al., 2021a) and MT-bench (Zheng et al., 2023). We conduct zero-shot evaluations on MMLU and report the average accuracy across all the subjects as $\mu_{\textrm{MU}}$. To better understand the model performance, we categorize the 57 subjects of MMLU into three groups: mathematics, medicine, and others, and also report the average accuracy of each group. Moreover, we report the average score across all eight skills of MT-bench as $\mu_{\textrm{MB}}$. The overall performance is reported as the average score of both $\mu_{\textrm{MU}}$ and $\mu_{\textrm{MB}}$, denoted as $\mu_{\textrm{BOTH}}$. Note that when computing $\mu_{\textrm{BOTH}}$, MT-bench scores are upscaled by 10×10$\times$10 × to range from 1 to 100, maintaining consistency with MMLU. More details are in Appendix D.
(a) MoS + CosSim + EMA
(b) MoS + Diff + EMA
@@ 6 Fine-tuning from Generalist to Specialist
We aim to fine-tune Llama-3-8B to specialize in mathematics using datasets from Section 4.1, referring to this modified version as MoSpec. For MoSpec with CosSim, we compute the cosine similarity between Mathematics and other datasets, including Mathematics itself. For MoSpec with Diff, we double the reward for the Mathematics dataset. For comparison, we fine-tune Llama-3-8B directly on Mathematics dataset in Section 4.1, denoted as MathLlama-3-8B. Both MoSpec and MathLlama-3-8B use the identical hyperparameters from Appendix C, except MathLlama-3-8B is fine-tuned for 12 epochs for fairness. We evaluate the models on math-related subjects in MMLU (0-shot, denoted as M-math), GSM8K (5-shot) (Cobbe et al., 2021), and MATH (5-shot) (Hendrycks et al., 2021b).
@@ 7 Related Work
The success of large language models (LLMs) heavily relies on their training datasets. Researchers gather or create extensive datasets (Raffel et al., 2020; Gao et al., 2021; Penedo et al., 2023; Wang et al., 2023; Li et al., 2023a; Cui et al., 2023; Wu et al., 2024b) Recent efforts focus on selecting data subsets to enhance training efficiency. Xie et al. (2023) estimate the quality of each subset in the pretraining dataset mixture using a small proxy model. Recent approaches filter out low-quality examples using perplexity (Li et al., 2023b; Marion et al., 2023).
@@ Appendix A Preliminary Study on Perplexity
Qwen1.5-0.5B Gemma-2B Llama-3-8B
$\text{PPL}_{\boldsymbol{\theta}_{0}}$ $\text{PPL}_{\boldsymbol{\theta}}$ $\Delta$ $\text{PPL}_{\boldsymbol{\theta}_{0}}$ $\text{PPL}_{\boldsymbol{\theta}}$ $\Delta$ $\text{PPL}_{\boldsymbol{\theta}_{0}}$ $\text{PPL}_{\boldsymbol{\theta}}$ $\Delta$
Mathematics 04.18 2.94 0.70 05.85 2.31 0.39 05.65 2.94 0.52
Medicine 08.38 4.45 0.53 08.60 2.95 0.34 05.86 2.51 0.43
General 06.05 4.01 0.66 09.66 3.51 0.36 04.25 2.51 0.59
NLP 37.70 7.98 0.21 49.28 4.78 0.10 29.79 4.19 0.14
Table 5: Preliminary results on perplexity. $\text{PPL}_{\boldsymbol{\theta}_{0}}$ and $\text{PPL}_{\boldsymbol{\theta}}$ are the average perplexity scores on each subset given by the original LLM backbone and the fine-tuned LLM with Prop. ($\tau=1$), respectively. $\Delta=\frac{\text{PPL}_{\boldsymbol{\theta}}}{\text{PPL}_{\boldsymbol{\theta}_{0}}}$ indicates the relative decrease in perplexity. A high value of $\Delta$ indicates the dataset is difficult to learn, while a lower value indicates the opposite.
@@
#exam. #words Inst.L Resp.L Turns
Mathematics 026.2K 03.4M 047.6 084.0 1.0
Medicine 005.2K 01.0M 036.5 147.4 1.0
General 009.3K 09.3M 054.3 202.4 3.6
NLP 062.6K 08.6M 127.9 009.7 1.0
Total 103.4K 22.3M 088.4 084.3 1.2
Appendix B Training Datasets
Appendix C Optimization
@@
Appendix D Evaluation
  • MMLU: Hendrycks et al. (2021a) propose the MMLU benchmark, covering 57 subjects across STEM, humanities, social sciences, and more. We categorize the subjects into three groups: mathematics, medicine, and others, and conduct zero-shot evaluations. We report the average accuracy for each group and the overall accuracy across all subjects, denoted as $\mu_{\textrm{MU}}$. Detailed subject categorization is available in Table 7.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated their extraordinary capabilities and are expected to proficiently master a diverse range of skills (Ouyang et al., 2022; Sanh et al., 2022; OpenAI, 2023; Anil et al., 2023b; Touvron et al., 2023a, b; Anil et al., 2023a; Mesnard et al., 2024), such as writing, reasoning, chatting, coding, and more, through supervised fine-tuning (SFT) and reinforcement learning from human feedback (RLHF) on an extensive collection of datasets from various sources (Bai et al., 2022; Longpre et al., 2023; Ding et al., 2023; Wang et al., 2024b). Each dataset contributes unique elements to the model’s skill set, but this diversity also brings challenges.
@@ 2 Preliminaries
Wang et al. (2020a) propose a general framework that automatically re-weighs training instances to enhance model performance, utilizing a validation set $D_{\textrm{dev}}$. This framework consists of two components: the model $\boldsymbol{\theta}$ and the scorer network $\boldsymbol{\psi}$. The scorer network $\boldsymbol{\psi}$ is designed to calculate a sampling probability for each training instance, which reflects its impact on validation performance. Training instances that have a greater similarity with $D_{\textrm{dev}}$ are allocated a higher probability, thus increasing their likelihood of being selected for model $\boldsymbol{\theta}$ updates. MoS is inspired by DDS but has key differences. Firstly, MoS focuses on rebalancing datasets, unlike DDS, which reweighs training instances. Secondly, MoS does not require prior knowledge of downstream applications, whereas DDS relies on validation set feedback, risking overfitting to that specific validation set. Thirdly, DDS uses the same architecture for the scorer network $\boldsymbol{\psi}$ and model $\boldsymbol{\theta}$, limiting its scalability, while MoS uses a simple MLP model as its scorer network (see Section 3.1).
@@ 3.2.3 Learning Trajectory
4 Experiments
4.1 Experimental Setup
#exam. #words Inst.L Resp.L Turns
Mathematics 026.2K 03.4M 047.6 084.0 1.0
Medicine 005.2K 01.0M 036.5 147.4 1.0
General 009.3K 09.3M 054.3 202.4 3.6
NLP 062.6K 08.6M 127.9 009.7 1.0
Total 103.4K 22.3M 088.4 084.3 1.2
Datasets
@@ 4.1 Experimental Setup
We apply MoS to three diverse model backbones, including Qwen1.5-0.5B (Bai et al., 2023), Gemma-2B (Mesnard et al., 2024), and Llama-3-8B (AI@Meta, 2024).
Optimization
@@
We compare MoS with several heuristic and dynamic baselines as follows:
  • Heuristic: Based on Equation 3, we consider proportional sampling (Prop., $\tau=1$), temperature sampling (Temp., $\tau=10$), and uniform sampling (Uni., $\tau=\infty$).
  • Dynamic: MultiDDS (Wang et al., 2020b) and MultiUAT (Wu et al., 2021) dynamically balance the dataset distribution using the gradient cosine similarity and model uncertainty on the validation sets as rewards, respectively. We sample 1K examples from each dataset as the validation sets for both approaches.
We do not include maximum cap and other instance selection methods as baselines because they do not fully utilize all available data.
@@
  • MMLU: Hendrycks et al. (2021a) propose the MMLU benchmark, covering 57 subjects across STEM, humanities, social sciences, and more. We categorize the subjects into three groups: mathematics, medicine, and others, and conduct zero-shot evaluations. We report the average accuracy for each group and the accuracy across all subjects, denoted as $\mu_{\textrm{MU}}$. Detailed subject categorization is in Appendix B.
@@ 6 Fine-tuning from Generalist to Specialist
We aim to fine-tune Llama-3-8B to specialize in mathematics using datasets from Section 4.1, referring to this modified version as MoSpec. For MoSpec with CosSim, we compute the cosine similarity between Mathematics and other datasets, including Mathematics itself. For MoSpec with Diff, we double the reward for the Mathematics dataset. For comparison, we fine-tune Llama-3-8B directly on Mathematics dataset in Section 4.1, denoted as MathLlama-3-8B. Both MoSpec and MathLlama-3-8B use the identical hyperparameters from Section 4.1, except MathLlama-3-8B is fine-tuned for 12 epochs for fairness. We evaluate the models on math-related subjects in MMLU (0-shot, denoted as M-math), GSM8K (5-shot) (Cobbe et al., 2021), and MATH (5-shot) (Hendrycks et al., 2021b).
@@ 7 Related Work
The success of large language models (LLMs) heavily relies on their training datasets. Researchers gather or create extensive datasets (Raffel et al., 2020; Gao et al., 2021; Penedo et al., 2023; Wang et al., 2023; Li et al., 2023a; Cui et al., 2023; Wu et al., 2024b; Wang et al., 2024a). Recent efforts focus on selecting data subsets to enhance training efficiency. Xie et al. (2023) estimate the quality of each subset in the pretraining dataset mixture using a small proxy model. Recent approaches filter out low-quality examples using perplexity (Li et al., 2023b; Marion et al., 2023).
@@ 9 Limitations
Acknowledgment
This work is partly supported by the ARC Future Fellowship FT190100039. This material is based on research partially supported by the DARPA Assured Neuro Symbolic Learning and Reasoning (ANSR) program under award number FA8750-23-2-1016.
@@ Appendix A Preliminary Study on Perplexity
Appendix B MMLU Subject Categorization
We present the detailed subject categorization of MMLU in Table 7.